# Panacea [Uncaptioned image]: Mitigating Harmful Fine-tuning for Large Language Models via Post-fine-tuning Perturbation

### v1

@@ Abstract
Harmful fine-tuning attack introduces significant security risks to the fine-tuning services. Main-stream defenses aim to vaccinate the model such that the later harmful fine-tuning attack is less effective. However, our evaluation results show that such defenses are fragile– with a few fine-tuning steps, the model still can learn the harmful knowledge. To this end, we do further experiment and find that an embarrassingly simple solution– adding purely random perturbations to the fine-tuned model, can recover the model from harmful behaviors, though it leads to a degradation in the model’s fine-tuning performance. To address the degradation of fine-tuning performance, we further propose Panacea, which optimizes an adaptive perturbation that will be applied to the model after fine-tuning. Panacea maintains model’s safety alignment performance without compromising downstream fine-tuning performance. Comprehensive experiments are conducted on different harmful ratios, fine-tuning tasks and mainstream LLMs, where the average harmful scores are reduced by up-to 21.5%, while maintaining fine-tuning performance. As a by-product, we analyze the adaptive perturbation and show that different layers in various LLMs have distinct safety coefficients. Source code available at https://github.com/w-yibo/Panacea.
Machine Learning, ICML
1 Tsinghua University; 2 Shenzhen Campus of Sun Yat-sen University; 3 Didichuxing Co. Ltd; 4 Nanyang Technological University. Correspondence to: Li Shen <mathshenli@gmail.com>.
@@ 1 Introduction
Fine-tuning-as-a-service (OpenAI, ) is a popular business service to enhance model’s performance for customized datasets, domain-specific tasks, and private needs. However, recent studies (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023; Yi et al., 2024a; Li et al., 2024a; Ye et al., 2024) identify a safety issue, the harmful fine-tuning attack (Figure 1), where the model’s safety alignment is compromised when the fine-tuning dataset contains harmful data, even a small amount of harmful data can introduce significant security vulnerabilities. Moreover, harmful fine-tuning is often unintentional, as datasets may contain latent unsafe data that is difficult for users to detect. Since service providers are responsible for the harmful outputs generated by the model, there is a clear need for more effective solutions.
Figure 1: The harmful fine-tuning attack for fine-tuning-as-a-service scenarios. In the alignment stage, pretrained LLMs are aligned using alignment data to produce aligned LLMs. In the fine-tuning stage, aligned LLMs are further fine-tuned using fine-tuning data, which may contain harmful data, potentially leading to unsafe fine-tuned models.
Vaccine (Huang et al., 2024e) and RepNoise (Rosati et al., 2024a) are two representative defenses against the harmful fine-tuning attack. Vaccine improves the aligned model’s resistance towards harmful knowledge learning by solving a mini-max problem, while RepNoise aims to erase the harmful knowledge of the aligned model by perturbing the hidden representation given harmful data. However, our evaluation results show that with more fine-tuning steps, the vaccinated model produced by the two methods are still suffering from the negative effect of harmful fine-tuning attack – their harmful score are significantly increased, though with a slower rate compared to the baseline without defense.
@@
How to add post-fine-tuning perturbation to the fine-tuned model, such that it can be recovered from harmful behavior without hurting downstream task performance much?
Driven by this question, we propose Panacea (Figure 2), an iterative gradient-based method to iterately search for the post-fine-tuning perturbation. Panacea aims to solve a max-maximize optimization problem, such that the added perturbation can maximally increase the model’s harmful loss, ensuring that it can effectively recover the model from harmful behaviors. The experiment results show that for different harmful ratios during fine-tuning, our method’s average harmful score is reduced by 10.4% up to 21.5%, while fine-tuning performance improves by 0.3% over the standard alignment method. The ablation study on the adaptive perturbation show that it can reduce harmful scores by up to 23.2%, while maintaining competitive fine-tuning performance. The visualization experiments also reveal different layers in various LLMs have distinct safety coefficients, consistent with previous findings and providing additional evidence for layer-wise safety research.
The main contributions of this work contain three parts:
1. We find that adding purely random perturbations to the fine-tuned models could recover the model from harmful behavior, but does cause a loss of fine-tuning performance.
2. To mitigate harmful fine-tuning while maintaining the fine-tuning performance, we propose Panacea, a post-fine-tuning solution that formulates a max-max optimization problem, where the optimized perturbation maximally increases the harmful loss.
3. Experiments evaluate Panacea across diverse settings, demonstrating its effectiveness and generalization, while visualizations reveal safety coefficients across LLM layers.
@@ 2 Related Work
Safety Alignment. The safety alignment requires the model to output content that is both helpful and harmless, and to be able to output a refusal answer when given harmful instructions. Existing methods typically rely on supervised fine-tuning (SFT), RLHF (Ouyang et al., 2022), and variations of RLHF (e.g., PPO, DPO) (Bai et al., 2022; Wu et al., 2023; Dong et al., 2023; Rafailov et al., 2023; Yuan et al., 2023; Huang et al., 2021). These methods construct a safety-aligned dataset, and recent approaches focus on enhancing and better utilizing the aligned dataset (Liu et al., 2023a, b; Ye et al., 2023; Tekin et al., 2024).
Figure 2: Post-fine-tuning perturbation. The fine-tuned model exhibits a high harmful score (HS:$\downarrow$). Adding random perturbation reduces the harmful score but also decreases fine-tuning accuracy (FA:$\uparrow$). In contrast, incorporating our post-fine-tuning perturbation (see Algorithm 1) effectively lowers the harmful score while maintaining fine-tuning performance.
@@
Harmful Fine-tuning. Fine-tuning-as-a-service becomes a mainstream method for LLMs API providers. Recent studies (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023; Li et al., 2024a; Ye et al., 2024; Lermen et al., 2023; He et al., 2024; Halawi et al., 2024; Chen et al., 2024; Hawkins et al., 2024; Poppi et al., 2024) show that LLMs trained with safety alignment can be jail-broken when fine-tuned on a dataset with a small amount of harmful data. In such cases, the model fails to refuse harmful instructions and outputs harmful responses. Many works (Rosati et al., 2024c; Leong et al., 2024; Anonymous, 2025c; Guo et al., 2024; Qi et al., 2024b; Rosati et al., 2024a) focus on analyzing the mechanisms of different harmful fine-tuning attacks. (Peng et al., 2024) proposes new safety metrics to evaluate harmful fine-tuning risk and (Rosati et al., 2024a) explores the safety risks when learning with reinforcement learning. Existing defenses can be categorized into three main categories (Huang et al., 2024c), i) Alignment stage solutions (Huang et al., 2024e; Rosati et al., 2024b; Tamirisa et al., 2024; Liu et al., 2024c; Pan et al., ; Huang et al., 2024b; Anonymous, 2025a; Liu et al., 2024a), ii) Fine-tuning stage solutions (Mukhoti et al., 2023; Bianchi et al., 2023; Zong et al., 2024; Wei et al., 2024; Huang et al., 2024d; Wang et al., 2024; Lyu et al., 2024; Eiras et al., 2024; Zhang et al., 2024; Li et al., 2024b; Shen et al., 2024; Anonymous, 2025b; Li et al., 2025; Du et al., 2024a; Choi et al., 2024; Luo et al., 2024; Qi et al., 2024a), iii) Post-fine-tuning stage solutions (Casper et al., 2024; Yi et al., 2024c; Du et al., 2024b; Hsu et al., 2024; Huang et al., 2024a; Zhu et al., 2024; Liu et al., 2024b; Wu et al., 2024; Yi et al., 2024b). The proposed method in this paper is applied in the post-fine-tuning stage, aiming to restore safety alignment without sacrificing fine-tuning performance.
@@
Mainstream defenses focusing on the alignment stage lack sufficient durability against harmful fine-tuning (Qi et al., 2024b), motivating exploration of the post-fine-tuning stage. Existing post-fine-tuning solutions typically add perturbations based on prior knowledge, such as a safety subspace (Hsu et al., 2024) or safety-critical parameters (Zhu et al., 2024; Yi et al., 2024b). In contrast, Panacea optimizes an adaptive perturbation during fine-tuning without relying on prior knowledge.
3 Revisiting Alignment Defenses
@@ 3.1 Problem Setup
Harmful Fine-tuning. Harmful fine-tuning poses a significant threat to LLMs service providers (OpenAI, ). The scenario is illustrated in Figure 1, where LLMs service providers use an alignment dataset to perform safety alignment on a pretrained model, transforming it into an aligned model. Users then upload a fine-tuning dataset containing harmful data to the service provider. The fine-tuned dataset is deployed on the service provider’s server and used to generate personalized outputs for the users.
Threat Models. Following  (Qi et al., 2023; Rosati et al., 2024b; Huang et al., 2024b), we assume that, during the fine-tuning stage, $p$ (percentage) of the fine-tuning dataset consists of harmful data (i.e., harmful instruction-harmful response pairs), while the $1-p$ of data represents benign fine-tuning data (e.g., math question-answer pairs (Cobbe et al., 2021)). Furthermore, we assume that harmful and benign data are inseparable within the fine-tuning dataset.
Defense Assumptions. Assume that LLM providers host an alignment dataset (harmful instruction-harmless response pairs) used during the alignment stage. Such a dataset is also assumed to be available by Vaccine (Huang et al., 2024e), RepNoise (Rosati et al., 2024b), BEA(Wang et al., 2024). Additionally, we assume availability of a harmful dataset (harmful instruction-harmful response pairs). This harmful dataset is also assumed to be available by existing methods, e.g., RepNoise (Rosati et al., 2024b), TAR (Tamirisa et al., 2024), Booster (Huang et al., 2024b). Both the alignment dataset and the harmful dataset can be obtained from existing open-sourced datasets (e.g., BeaverTails).
3.2 Revisiting Alignment Defenses
In this section, we explore the existing alignment stage solutions during the harmful fine-tuning and discuss the feasibility of introducing random perturbation in the post-fine-tuning stage, thus inspiring our proposed method.
@@ 3.2 Revisiting Alignment Defenses
We select the SFT method (vanilla supervised fine-tuning with the alignment dataset), Vaccine (Huang et al., 2024e), and RepNoise (Rosati et al., 2024b), and perform fine-tuning on a fine-tuning dataset containing a small proportion (0.1) of harmful samples. As shown in Figure 3, the three methods start with different levels of harmful loss; however, as training progresses, they all achieve lower harmful loss, which corresponds to an increase in harmful score, making the model harmful. This naturally leads us to consider:
Can simply adding a random perturbation after the fine-tuning to increase the harmful loss to restore the safety alignment?
Figure 4: Model statistics (Left: harmful loss, Right: harmful $\downarrow$ and fine-tuning $\uparrow$) for fine-tuned model with noise intensities of 0 (no noise), 0.001, 0.01, 0.05, 0.1. (HS of 0.1 is greater than 80, FA of 0.1 is 0.7, and is not shown.)
To evaluate it, we design the experiments that add Gaussian noise with intensities of 0.001, 0.01, 0.05, and 0.1 to the fine-tuned model. The experimental results shown in Figure 4 indicate that adding random Gaussian noise increases harmful loss, demonstrating that random perturbations have the potential to prevent the harmful fine-tuning. We further measure the effects of adding random perturbations. As shown in Figure 4, the quantitative results reveal that random perturbations reduce the model’s harmful score. And the reduction effect improves as the noise intensity increases. However, excessive noise (0.1 of intensity) ultimately degrades the model. At the same time, random perturbations significantly impair the fine-tuned model’s performance.
Derived Insight. As the harmful dataset has the potential to direct the perturbation towards altering the harmful loss, thereby reducing perturbation’s impact on fine-tuning performance, we could optimize an perturbation with harmful dataset during fine-tuning stage, which recover the safety alignment while maintaining the fine-tuning performance.
@@ 4 Methodology
At the fine-tuning stage, we aim to optimize an adaptive perturbation that maximally increases the harmful loss. This perturbation is then added to the fine-tuned model after the fine-tuning process. Therefore, our method is a post-fine-tuning stage solution, formulated as a max-maximize optimization, as follows:
@@
As shown in Algorithm 1, the optimization process consists of four key steps: First, a batch of fine-tuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$ is used to compute the gradient ~$\widetilde{\nabla}g(\boldsymbol{w}_{t})$, where $\widetilde{\nabla}$ notes a batch of gradient. Second, a batch of harmful data $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$ is sampled to compute the harmful gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t})$. Third, the perturbation is computed (Eq. 2) and applied to update the harmful gradient, yielding ~$\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$. Lastly, the combined gradient ~$\widetilde{\nabla}f(\boldsymbol{w}_{t})$ is computed and used to update the model parameters, with the final perturbation applied to obtain the re-aligned parameters $\boldsymbol{w}_{\varepsilon}$.
@@
2:    Sample a batch of finetuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
@@
4:    Compute gradient ~$\widetilde{\nabla}g(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
5:    Compute gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
@@
7:    Compute gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
8:    ~$\widetilde{\nabla}f(\boldsymbol{w}_{t})=\lambda(\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})-\widetilde{\nabla}h(\boldsymbol{w}_{t}))-\widetilde{\nabla}g(\boldsymbol{w}_{t})$
@@
Our proposed algorithm, dubbed as Panacea, is named after an alignment-stage defense Vaccine (Huang et al., 2024e). However, we note that these two defenses are fundamentally different. Vaccine vaccinates the LLM to enhance its robustness at the alignment stage in order to counter attacks launched after alignment. By constrast, our algorithm Panacea belongs to the post-fine-tuning stage, where it introduces an optimized adaptive perturbation to restore the model’s safety alignment after the fine-tuning stage. We also show in experiment that Panacea outperforms Vaccine in both the two concerned metrics, i.e., harmful score and fine-tuning accuracy.
Table 1: Performance comparison of different harm ratio.
Method Harmful Score($\downarrow$) Finetuning Accuracy($\uparrow$)
Clean 0.05 0.1 0.15 0.2 0.25 Avg. Clean 0.05 0.1 0.15 0.2 0.25 Avg.
SFT 5.2 27.5 45.8 56.2 67.0 69.4 45.2 16.4 16.0 16.2 15.4 15.2 16.8 16.0
Vaccine 2.3 15.6 25.4 40.3 55.2 65.9 34.1 14.2 13.9 13.5 13.0 13.6 13.4 13.6
RepNoise 2.7 22.2 32.0 42.9 54.0 62.4 36.0 15.7 16.1 15.8 14.9 14.0 14.8 15.2
Booster 4.6 21.0 42.3 60.7 69.3 72.5 45.1 17.8 17.8 17.6 17.1 16.2 14.7 16.9
Panacea 1.8 9.9 20.1 29.1 34.8 46.3 23.7 15.0 16.3 16.7 17.0 16.2 17.4 16.3

### v2

# Panacea: Mitigating Harmful Fine-tuning for Large Language Models via Post-fine-tuning Perturbation

@@ Abstract
Harmful fine-tuning attack introduces significant security risks to the fine-tuning services. Main-stream defenses aim to vaccinate the model such that the later harmful fine-tuning attack is less effective. However, our evaluation results show that such defenses are fragile– with a few fine-tuning steps, the model still can learn the harmful knowledge. To this end, we do further experiment and find that an embarrassingly simple solution– adding purely random perturbations to the fine-tuned model, can recover the model from harmful behaviors, though it leads to a degradation in the model’s fine-tuning performance. To address the degradation of fine-tuning performance, we further propose Panacea, which optimizes an adaptive perturbation that will be applied to the model after fine-tuning. Panacea maintains model’s safety alignment performance without compromising downstream fine-tuning performance. Comprehensive experiments are conducted on different harmful ratios, fine-tuning tasks and mainstream LLMs, where the average harmful scores are reduced by up-to 21.2%, while maintaining fine-tuning performance. As a by-product, we analyze the adaptive perturbation and show that different layers in various LLMs have distinct safety affinity, which coincide with finding from several previous study. Source code available at https://github.com/w-yibo/Panacea.
@@ 1 Introduction
Figure 1: The harmful fine-tuning attack for fine-tuning-as-a-service scenarios. Pretrained LLMs are aligned using alignment data to produce aligned LLMs. Aligned LLMs are further fine-tuned using fine-tuning data that may contain harmful data, leading to unsafe fine-tuned models.
Fine-tuning-as-a-service OpenAI is a popular business service to enhance model’s performance for customized datasets, domain-specific tasks, and private needs. However, recent studies Qi et al. (2023); Zhan et al. (2023); Yang et al. (2023); Yi et al. (2024a); Li et al. (2024b); Ye et al. (2024) identify a safety issue, the harmful fine-tuning attack (Figure 1), where the model’s safety alignment is compromised when the fine-tuning dataset contains harmful data, even a small amount of harmful data can introduce significant security vulnerabilities. Moreover, harmful fine-tuning is often unintentional, as datasets may contain latent unsafe data that is difficult for users to detect. Since service providers are responsible for the harmful outputs generated by the model, there is a clear need for effective solutions.
Vaccine Huang et al. (2024e) and RepNoise Rosati et al. (2024b) are two representative defenses against the harmful fine-tuning attack. Vaccine improves the aligned model’s resistance towards harmful knowledge learning by solving a mini-max problem, while RepNoise aims to erase the harmful knowledge of the aligned model by perturbing the hidden representation given harmful data. However, our evaluation results show that with more fine-tuning steps, the vaccinated model produced by the two methods are still suffering from the negative effect of harmful fine-tuning attack – their harmful score are significantly increased, though with a slower rate compared to the baseline without defense.
@@
How to add post-fine-tuning perturbation to the fine-tuned model, such that it can be recovered to safe state without hurting downstream performance too much?
[…]
Driven by this question, we propose Panacea (Figure 2), an iterative gradient-based method to iterately search for the post-fine-tuning perturbation. Panacea aims to solve a max-maximize optimization problem, such that the added perturbation can maximally increase the model’s harmful loss, ensuring that it can effectively recover the model from harmful behaviors. The experiment results show that for different harmful ratios during fine-tuning, our method’s average harmful score is reduced by up to 21.2%, while fine-tuning performance improves by 0.4% over the standard alignment method. The ablation study on the adaptive perturbation show that it can reduce harmful scores by up to 23.2%, while maintaining competitive fine-tuning performance. The visualization experiments also reveal different layers in various LLMs have distinct safety coefficients, consistent with previous findings and providing additional evidence for layer-wise safety research.
The main contributions of this paper: i) We find that adding purely random perturbations to the fine-tuned models could recover the model from harmful behavior, but does cause a loss of fine-tuning performance. ii) To mitigate harmful fine-tuning while maintaining the fine-tuning performance, we propose Panacea, a post-fine-tuning solution that formulates a max-max optimization problem, where the optimized perturbation maximally increases the harmful loss. iii) Experiments evaluate Panacea across diverse settings, demonstrating its effectiveness and generalization, while visualizations reveal safety coefficients across LLM layers.
@@ 2 Related Work
Safety Alignment. The safety alignment requires the model to output content that is both helpful and harmless, and to be able to output a refusal answer when given harmful instructions. Existing methods typically rely on supervised fine-tuning (SFT), RLHF Ouyang et al. (2022), and variations of RLHF (e.g., PPO, DPO) Bai et al. (2022); Wu et al. (2023); Dong et al. (2023); Rafailov et al. (2023); Yuan et al. (2023); Huang et al. (2021). These methods construct a safety-aligned dataset, and recent approaches focus on enhancing and better utilizing the aligned dataset Liu et al. (2023a, b); Ye et al. (2023); Tekin et al. (2024); Huang et al. (2024f).
@@
Harmful Fine-tuning. Fine-tuning-as-a-service becomes a mainstream method for LLMs API providers. Recent studies Qi et al. (2023); Zhan et al. (2023); Yang et al. (2023); Li et al. (2024b); Ye et al. (2024); Lermen et al. (2023); He et al. (2024); Halawi et al. (2024); Chen et al. (2024); Hawkins et al. (2024); Poppi et al. (2024); Guan et al. (2025); Djuhera et al. (2025a); Pandey et al. (2025); Gloaguen et al. (2025); Murphy et al. (2025); Wallace et al. (2025); Hahm et al. (2025); Shao et al. (2025); Li et al. (2025d); Huang et al. (2025); Xu et al. (2025); Davies et al. (2025); Kazdan et al. (2025); Betley et al. (2025); Upadhayay and Behzadan (2025); Anonymous (2025g, b) show that LLMs trained with safety alignment can be jail-broken when fine-tuned on a dataset with a small amount of harmful data. In such cases, the model fails to refuse harmful instructions and outputs harmful responses. Many works Rosati et al. (2024a, c); Leong et al. (2024); Hsiung et al. ; Guo et al. (2024); Qi et al. (2024b); Li et al. (2025a); Lee et al. (2025); O’Brien et al. (2025); Che et al. (2025); Chen et al. (2025c); Ponkshe et al. (2025); Uppal and Tankala (2025); kaczér2025intrainingdefensesemergentmisalignment; Hossain et al. (2025); Anonymous (2025f) focus on analyzing the mechanisms of different harmful fine-tuning attacks. Peng et al. (2024) proposes new safety metrics to evaluate harmful fine-tuning risk and Rosati et al. (2024a) explores the safety risks when learning with reinforcement learning. Existing defenses can be categorized into three main categories Huang et al. (2024c), i) Alignment stage solutions Huang et al. (2024e); Rosati et al. (2024b); Tamirisa et al. (2024); Liu et al. (2024c); Pan et al. ; Huang et al. (2024b); Zhao et al. (2025b); Liu et al. (2024a); Cheng et al. (2025); Fan et al. (2025); Cao (2025); Wang et al. (2025c); Yi et al. (2025a); Zheng et al. (2025); Wang et al. (2025a); Chen et al. (2025b); Rosati et al. (2025); Yang et al. (2024b); Perin et al. (2025); Feng et al. (2025); Anonymous (2025a), ii) Fine-tuning stage solutions Mukhoti et al. (2023); Bianchi et al. (2023); Zong et al. (2024); Wei et al. (2024); Huang et al. (2024d); Wang et al. (2024); Lyu et al. (2024); Eiras et al. (2024); Zhang et al. (2024); Li et al. (2024a); Shen et al. (2024); Li and Kim (2024); Li et al. (2025c); Du et al. (2024a); Choi et al. (2024); Luo et al. (2024); Qi et al. (2024a); Hu et al. (2025); Peng et al. (2025); Zhao et al. (2025a); Fu and Barez (2025); Liu et al. (2025); Li (2025); Wang et al. (2025b); Wu et al. (2025); Chen et al. (2025a); Luo et al. (2025); Xiao et al. (2025); Ham et al. (2025); Yang et al. (2025b); Li et al. (2025b); Das et al. (2025); Kim et al. (2025b); Yi et al. (2025b); Youstra et al. (2025); Du et al. (2025a); Kim et al. (2025a); Anonymous (2025c, d); Zhang et al. (2025), iii) Post-fine-tuning stage solutions Casper et al. (2024); Yi et al. (2024c); Du et al. (2024b); Hsu et al. (2024); Huang et al. (2024a); Zhu et al. (2024); Liu et al. (2024b); Wu et al. (2024); Yi et al. (2024b); Gong et al. (2025); Djuhera et al. (2025b); Yang et al. (2025a); Lu et al. (2025); Ao et al. (2025); Zhou et al. (2025); Han et al. (2025); Du et al. (2025b); Anonymous (2025e); Jiang et al. (2025). The proposed method in this paper is applied in the post-fine-tuning stage, aiming to restore safety alignment without sacrificing fine-tuning performance.
@@
Mainstream defenses focusing on the alignment stage lack sufficient durability against harmful fine-tuning Qi et al. (2024b), motivating exploration of the post-fine-tuning stage. Existing post-fine-tuning solutions typically add perturbations based on prior knowledge, such as a safety subspace Hsu et al. (2024) or safety-critical parameters Zhu et al. (2024); Yi et al. (2024b). In contrast, Panacea optimizes an adaptive perturbation during fine-tuning without relying on prior knowledge.
3 Preliminaries
@@ 3.1 Problem Setup
Harmful Fine-tuning. Harmful fine-tuning poses a significant threat to LLMs service providers OpenAI . The scenario is illustrated in Figure 1, where LLMs service providers use an alignment dataset to perform safety alignment on a pretrained model, transforming it into an aligned model. Users then upload a fine-tuning dataset containing harmful data to the service provider. The fine-tuned dataset is deployed on the service provider’s server and used to generate personalized outputs for the users.
Threat Models. Following  Qi et al. (2023); Rosati et al. (2024b); Huang et al. (2024b), we assume that, during the fine-tuning stage, $p$ (percentage) of the fine-tuning dataset consists of harmful data (i.e., harmful instruction-harmful response pairs), while the 1p1-p of data represents benign fine-tuning data (e.g., math question-answer pairs Cobbe et al. (2021)). Furthermore, we assume that harmful and benign data are inseparable within the fine-tuning dataset.
Defense Assumptions. Assume that LLM providers host an alignment dataset (harmful instruction-harmless response pairs) used during the alignment stage. Such a dataset is also assumed to be available by Vaccine Huang et al. (2024e), RepNoise Rosati et al. (2024b), BEAWang et al. (2024). Additionally, we assume availability of a harmful dataset (harmful instruction-harmful response pairs). This harmful dataset is also assumed to be available by existing methods, e.g., RepNoise Rosati et al. (2024b), TAR Tamirisa et al. (2024), Booster Huang et al. (2024b). Both the alignment dataset and the harmful dataset can be obtained from existing open-sourced datasets (e.g., BeaverTails).
3.2 Exploration Study
We first explore the existing alignment stage solutions against harmful fine-tuning and show by statistical results that these designs still cannot eliminate the risk of harmful fine-tuning.
@@ 3.2 Exploration Study
Pre-fine-tuning defenses lack robustness. We select the SFT method (vanilla supervised fine-tuning with the alignment dataset), and two pre-fine-tuning defenses Vaccine Huang et al. (2024e) and RepNoise Rosati et al. (2024b) as baseline for evaluation. We perform fine-tuning on a fine-tuning dataset containing a small proportion (0.1) of harmful samples. As shown in Figure 3, the three methods start with different levels of harmful loss; however, as training progresses, they all achieve lower harmful loss, which corresponds to an increase in harmful score, making the model harmful. Fundamentally, it seems that pre-fine-tuning defense is not the best direction to counter fine-tuning attack as the fine-tuning attack can still effectively subvert the model’s safety alignment with more fine-tuning steps.
Exploring post-fine-tuning defense. Given that the pre-fine-tuning procedure cannot effective resist the attack, this naturally leads us to consider a potential defense baseline to counter the attack after the attack has been implanted to the model. Our initial idea is simple– we want to test whether a random perturbation over the model weights can restore the model from its harmful state. Specifically, the following question needs to be explored:
Can simply add a random perturbation after the fine-tuning to increase the harmful loss to restore the safety alignment?
Random perturbation recovers model to a safety state, but it hurts model’s performance. We design the experiments that add Gaussian noise with intensities of 0.001, 0.01, 0.05, and 0.1 to the fine-tuned model. The experimental results shown in Figure 4 indicate that adding random Gaussian noise increases harmful loss, demonstrating that random perturbations have the potential to prevent the harmful fine-tuning. We further measure the effects of adding random perturbations. As shown in Figure 4, the quantitative results reveal that random perturbations reduce the model’s harmful score. And the reduction effect improves as the noise intensity increases. However, as shown in the right of Figure 4, random perturbations significantly impair the fine-tuned model’s performance as the fine-tuning accuracy is also significantly downgraded with the increase of noise intensity.
[…]
We need a more carefully crafted post-fine-tuning perturbation. As post-fine-tuning perturbation has potential to recover the model to a safe state but it comes with a degradation of model’s generalization performance, we need to explore a better way to craft such perturbation. We will discuss this better perturbation crafting method in the following section.
@@ 4 Methodology
In this section, we discuss our method to craft a better post-fine-tuning perturbation to recover the model from fine-tuning attack. To search for such perturbation, we operate an extra optimization at the fine-tuning stage. Specifically, at the fine-tuning stage, we aim to optimize an adaptive perturbation that maximally increases the harmful loss. This perturbation is then added to the fine-tuned model after the fine-tuning process. Formally, our method can be formulated as a max-maximize optimization, as follows:
@@
2:  Sample a batch of fine-tuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
@@
4:  Compute gradient $\widetilde{\nabla}g(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
5:  Compute gradient $\widetilde{\nabla}h(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
@@
7:  Compute gradient $\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
8:  $\widetilde{\nabla}f(\boldsymbol{w}_{t})=\lambda(\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})-\widetilde{\nabla}h(\boldsymbol{w}_{t}))-\widetilde{\nabla}g(\boldsymbol{w}_{t})$
@@
As shown in Algorithm 1, the optimization process consists of four key steps: First, a batch of fine-tuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$ is used to compute the gradient $\widetilde{\nabla}g(\boldsymbol{w}_{t})$, where $\widetilde{\nabla}$ notes a batch of gradient. Second, a batch of harmful data $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$ is sampled to compute the harmful gradient $\widetilde{\nabla}h(\boldsymbol{w}_{t})$. Third, the perturbation is computed (Eq. 2) and applied to update the harmful gradient, yielding $\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$. Lastly, the combined gradient $\widetilde{\nabla}f(\boldsymbol{w}_{t})$ is computed and used to update the model parameters, with the final perturbation applied to obtain the re-aligned parameters $\boldsymbol{w}_{\varepsilon}$.
Our proposed algorithm, dubbed as Panacea, is named after an alignment-stage defense Vaccine Huang et al. (2024e). However, we note that these two defenses are fundamentally different. Vaccine vaccinates the LLM to enhance its robustness at the alignment stage in order to counter attacks launched after alignment. By contrast, our algorithm Panacea belongs to the post-fine-tuning stage, where it introduces an optimized adaptive perturbation to restore the model’s safety alignment after the fine-tuning stage. Since Panacea does not require access to the alignment stage, it can be directly applied to already aligned LLMs such as Llama2-7B-Chat. We present experimental results on such aligned models in Table 4, demonstrating the broad applicability of our method. Furthermore, our experiments show that Panacea outperforms Vaccine on both key metrics—harmful score and fine-tuning accuracy.
Of note, when we prepare the camera ready of this paper, we find a previous work Security Vectors  Zhou et al. (2023) follow a similar idea with Panacea. Both Security Vectors and Panacea aim to make sure that the harmful loss can be sufficiently reduced during fine-tuning, and then add a perturbation after fine-tuning to remove the harmful knowledge. However, there is two difference between Security vector and Panacea. The first implementation difference is that security vector learns the perturbation before fine-tuning while Panacea learns the perturbation during fine-tuning. The second difference, which is the major difference is what the perturbation is and how it leads to increase of harmful loss. Panacea aim to find a perturbation that maximally increase the loss harmful loss and thereby adding the perturbation can sufficiently unlearn harmful knowledge. In contrast, the goal of security vector is to distill harmful knowledge into a harmful component before fine-tuning, and removing this component (which can be seen as a perturbation as well) to increase harmful loss. However, their formulation can not explicitly guarantee that the increase of harmful loss is maximized after adding perturbation (i.e., de-activating their security vector).